Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Help!! discrepancy in gmat prep & OG answer-700+ spoiled

Expert replies
by munaf » Sun Dec 20, 2009 10:11 pm
16) A basket has 5 apples, one is spoiled. If Henry picked 2 apples simultaneously and at random, what is probability that the 2 selected apples will include the spoiled one.
a. 1�5
b. 2�10
c. 2�5
d. 1�2
e. 3�5

gUYS THIS PROBLEM WAS DISCUSSED EARLIER AND THE OA IS D
I GOT THE SAME QUESTION ON GMAT PREP YESTERDAY AND I WAS BLOWN APART TO SEE THE RIGHT ANSWER AS C...

HELP!!!!! I HAVE TEST NEXT WEEK.
Join the discussion
Source: — Problem Solving |

by Perminology » Sun Dec 20, 2009 10:50 pm
I've seen this happen on multiple tests, but more so on Power Prep than GMAT Prep. Guess there isn't much you can do. Shrugs.
Join the discussion

by Ian Stewart » Sun Dec 20, 2009 11:08 pm
munaf wrote:16) A basket has 5 apples, one is spoiled. If Henry picked 2 apples simultaneously and at random, what is probability that the 2 selected apples will include the spoiled one.
a. 1�5
b. 2�10
c. 2�5
d. 1�2
e. 3�5

gUYS THIS PROBLEM WAS DISCUSSED EARLIER AND THE OA IS D
I GOT THE SAME QUESTION ON GMAT PREP YESTERDAY AND I WAS BLOWN APART TO SEE THE RIGHT ANSWER AS C...

HELP!!!!! I HAVE TEST NEXT WEEK.
Who is claiming the answer is 1/2? That isn't the correct answer. Imagine you line up the five apples, and you will pick the first two. There is a 2/5 chance that the spoiled apple is among the first two of the five apples, so the probability is 2/5.
For online GMAT math tutoring, or to buy my higher-level Quant books and problem sets, contact me at ianstewartgmat at gmail.com

ianstewartgmat.com
Join the discussion

by munaf » Sun Dec 20, 2009 11:10 pm
can you please explain how..asap
Join the discussion

by papgust » Sun Dec 20, 2009 11:17 pm
Munaf, we have already discussed this prob and the answer is indeed 2/5. We didn't claim that the answer is D.

https://www.beatthegmat.com/700-spoiled- ... 49344.html
Join the discussion

by munaf » Sun Dec 20, 2009 11:29 pm
Please see the two examples below and their solutions posted on this forum-majority of the responses were for the same answers. i.e for the 1st question majority explained their answer as D-1/2 and for the second question the answer as D-21/45.
I guess the solution for the 1st answer is wrong here-the numerator should be 1*4c1 and not 1+4c1.this is where the problem lies..guys please tell me the right way to approach these problems.

16) A basket has 5 apples, one is spoiled. If Henry picked 2 apples simultaneously and at random, what is probability that the 2 selected apples will include the spoiled one.
a. 1�5
b. 2�10
c. 2�5
d. 1�2
e. 3�5
Solution-

Here as per my understanding this question can be solved in this way:

Henry picks 2 apples from 5 apples in the basket so total possible ways is:5c2= 10.

Possibility that 2 selected apples include spoiled one is (Probality of getting one spolied one + Probability of getting one good one from the 4 apples)

this probability will be 1 + 4c1 = 5

Hence the probability of getting one spoiled among the two apples picked randomly is 5/10 = 1/2.

Hi, can someone help me with this problem- it seems simple, but I must be mistaken in my logic. Its from the Powerprep software Exam 2.

A certain basket contains 10 apples, 7 of which are red and 3 of which are green. If 3 different apples are to be selected at random from the basketbal, what is the probability that 2 of the apples selected will be red and 1 will be green?

A) 7/40
B) 7/20
C) 49/100
D) 21/40
E) 7/10

I thought it would just be 7/10 * 6/9 * 3/8, but thats wrong.
I'll post the answer after some discussion -- thanks.
Solution----
I think it's pretty simple...Let's C if I am right!

7C2 * 3C1 / 10C3

= 21/40

= Choice D

Cheers
Join the discussion

by munaf » Sun Dec 20, 2009 11:33 pm
sorry guys......the Gmat pressure is building as the test is approaching..i saw the explanations and rectified the mistake.
actuall i had seen the explanation for the spoiled apples question which was wrong!!!!

Extremely sorry
Join the discussion

by for120 » Mon Dec 21, 2009 5:01 am
Correct Solution

put question into 2 steps

if u pick one apple among X apple, 1/X chance get spoiled apples; 1-1/X chance get a good one.

if pick two apples, chance to get spoiled apple=1/5 (chance get that spoiled apple)*100%(whatever another apple is good or spoiled)+4/5 chance get a good one* 1/4(chance get that spoiled apple among the remaing four)=1/5*100%+4/5*1/4=2/5
Join the discussion

by Testluv » Mon Dec 21, 2009 7:12 am
munaf wrote:sorry guys......the Gmat pressure is building as the test is approaching..i saw the explanations and rectified the mistake.
actuall i had seen the explanation for the spoiled apples question which was wrong!!!!

Extremely sorry
Hey don't worry about it, mistakes happen. Just try to keep cool and calm in your final days before the test--that's important!
Kaplan Teacher in Toronto
Join the discussion